# The Starlink Robot: A Platform and Dataset for Mobile Satellite Communication

### v2

@@ Abstract.
The integration of satellite communication into mobile devices represents a paradigm shift in connectivity, yet the performance characteristics under motion and environmental occlusion remain poorly understood. We present the Starlink Robot, the first mobile robotic platform equipped with Starlink satellite internet, comprehensive sensor suite including upward-facing camera, LiDAR, and IMU, designed to systematically study satellite communication performance during movement. Our multi-modal dataset captures synchronized communication metrics, motion dynamics, sky visibility, and 3D environmental context across diverse scenarios including steady-state motion, variable speeds, and different occlusion conditions. This platform and dataset enable researchers to develop motion-aware communication protocols, predict connectivity disruptions, and optimize satellite communication for emerging mobile applications from smartphones to autonomous vehicles. The project is available at https://github.com/StarlinkRobot.
@@ 1. Introduction
Our contribution extends beyond the platform itself. We present a multi-modal dataset that synchronizes Starlink performance metrics – including latency, upload and download throughput, and signal quality indicators – with high-frequency motion data and environmental observations. This dataset captures diverse scenarios from steady locomotion to variable speeds, from open sky to heavily occluded urban environments. By releasing both our platform design and collected data, we provide the research community with tools to develop the next generation of mobile satellite communication systems.
@@ 2. Related Works
The rapid deployment of LEO satellite constellations has sparked significant research interest in characterizing their performance. The Starlink academic community, particularly through the University of Victoria’s PanLab, has produced comprehensive studies of Starlink’s several static performance characteristics (Zhao and Pan, 2024a; Yang et al., 2024; Zheng et al., 2024; Tian et al., 2024; Kamel et al., 2024; Ahangarpour et al., 2024; Li et al., 2025; Wang et al., 2024; Liu et al., 2025; Zheng et al., 2025; Hu et al., 2020; Yang et al., 2021; Deng et al., 2022; Wang et al., 2023; Pan et al., 2023, 2024; Zhao and Pan, 2024b). Kassem et al. analyzed throughput variations across geographic locations, revealing how latitude affects connection quality due to satellite density differences (Kassem et al., 2022). Muhammad et al. investigated weather impacts, demonstrating that rain fade affects Starlink less severely than traditional geostationary satellites due to shorter signal paths (Ullah et al., 2025). These foundational studies establish baseline performance metrics but explicitly acknowledge the limitation of stationary measurements.
@@ 3.2. Software Architecture
The Starlink data collection presents unique challenges as the terminal doesn’t expose a direct API. We employ a multi-layered approach: parsing the terminal’s gRPC status interface provides 1Hz updates on signal quality, obstruction detection, and aggregate throughput. For finer-grained analysis, we capture packet-level data using tcpdump on the terminal’s Ethernet interface, enabling microsecond-resolution latency measurements. Continuous active measurements complement passive monitoring – ICMP probes to geographically distributed servers run at 10Hz, while TCP and UDP throughput tests execute every 30 seconds to capture performance variations.
@@ 4. Dataset Description
Communication Performance Metrics. The core of our dataset is the Starlink communication measurements shown in Figure 8. We collect: (1) Terminal-reported statistics at 1 Hz including downlink/uplink throughput, RTT, SNR, and obstruction state; (2) Active network measurements with ICMP probes at 10 Hz to multiple servers; (3) TCP/UDP throughput tests every 30 seconds. All communication data is timestamped and synchronized with sensor data.

### v3

@@ Abstract.
The integration of satellite communication into mobile devices represents a paradigm shift in connectivity, yet the performance characteristics under motion and environmental occlusion remain poorly understood. We present the Starlink Robot, the first mobile robotic platform equipped with Starlink satellite internet, comprehensive sensor suite including upward-facing camera, LiDAR, and IMU, designed to systematically study satellite communication performance during movement. Our multi-modal dataset captures synchronized communication metrics, motion dynamics, sky visibility, and 3D environmental context across diverse scenarios including steady-state motion, variable speeds, and different occlusion conditions. This platform and dataset enable researchers to develop motion-aware communication protocols, predict connectivity disruptions, and optimize satellite communication for emerging mobile applications from smartphones to autonomous vehicles. In this work, we use LEOViz for real-time satellite tracking and data collection. The starlink robot project is available at https://github.com/StarlinkRobot.
@@ 1. Introduction
Our contribution extends beyond the platform itself. We present a multi-modal dataset that synchronizes Starlink performance metrics – including latency, upload and download throughput, and signal quality indicators – with high-frequency motion data and environmental observations. Satellite tracking data is collected using LEOViz (Ahangarpour et al., 2024; Zhao and Pan, 2025), which provides real-time visualization of satellite positions and connection status. This dataset captures diverse scenarios from steady locomotion to variable speeds, from open sky to heavily occluded urban environments. By releasing both our platform design and collected data, we provide the research community with tools to develop the next generation of mobile satellite communication systems.
@@ 2. Related Works
The rapid deployment of LEO satellite constellations has sparked significant research interest in characterizing their performance. The Starlink academic community, particularly through the University of Victoria’s PanLab, has produced comprehensive studies of Starlink’s several static performance characteristics (Zhao and Pan, 2024a; Yang et al., 2024; Zheng et al., 2024; Tian et al., 2024; Kamel et al., 2024; Ahangarpour et al., 2024; Li et al., 2025; Wang et al., 2024; Liu et al., 2025; Zheng et al., 2025; Hu et al., 2020; Yang et al., 2021; Deng et al., 2022; Wang et al., 2023; Pan et al., 2023, 2024; Zhao and Pan, 2024b). Among these, LEOViz (Ahangarpour et al., 2024; Zhao and Pan, 2025) provides real-time visualization and tracking of Starlink satellites, displaying satellite positions, elevation angles, and connection status. In our work, we utilize LEOViz as our data collection and visualization tool, running it alongside our robotic platform to record satellite tracking information during mobile experiments. The satellite data captured by LEOViz enables us to analyze satellite communication performance during movement. Kassem et al. analyzed throughput variations across geographic locations, revealing how latitude affects connection quality due to satellite density differences (Kassem et al., 2022). Muhammad et al. investigated weather impacts, demonstrating that rain fade affects Starlink less severely than traditional geostationary satellites due to shorter signal paths (Ullah et al., 2025). These foundational studies establish baseline performance metrics but explicitly acknowledge the limitation of stationary measurements.
@@ 3.2. Software Architecture
The Starlink data collection presents unique challenges as the terminal doesn’t expose a direct API. We employ a multi-layered approach: for satellite tracking and constellation data, we use LEOViz (Ahangarpour et al., 2024; Zhao and Pan, 2025), which handles the parsing of Starlink’s gRPC status interface and provides 1Hz updates on satellite positions, azimuth/elevation angles, signal quality, and connection status. LEOViz visualizes this data in real-time and we record its output during our experiments.
@@ 4. Dataset Description
Communication Performance Metrics. The core of our dataset is the Starlink communication measurements shown in Figure 8. We collect with LEOViz: (1) Terminal-reported statistics at 1 Hz including downlink/uplink throughput, RTT, SNR, and obstruction state; (2) Active network measurements with ICMP probes at 10 Hz to multiple servers; (3) TCP/UDP throughput tests every 30 seconds. All communication data is timestamped and synchronized with sensor data.
@@ 7. Conclusion
8. Acknowledgements
We thank the developers of LEOViz (Ahangarpour et al., 2024; Zhao and Pan, 2025) for providing the open-source satellite tracking and visualization tool that enabled the satellite data collection component of our research. LEOViz was instrumental in capturing real-time Starlink satellite positions, elevation angles, and connection status during our mobile experiments. We also acknowledge the open-source robotics community for the ROS framework and associated tools that facilitated our multi-modal data synchronization.
This work was supported by the UKRI Future Leaders Fellowship [MR/V025333/1] (RoboHike).